OpenStudy (anonymous):

use order of operation to solve. 25 divided by (8-3)^2 -1

OpenStudy (jennyrlz):

Hello :)

OpenStudy (jennyrlz):

Order of operations has a very meet trick i can teach you ;)

OpenStudy (jennyrlz):

Pemdas

OpenStudy (jennyrlz):

Parenthesis, exponents, multiply, divide, add, and subtract

OpenStudy (jennyrlz):

This signifies the order you should operate in

OpenStudy (jennyrlz):

I am on my ipad so I can't see if you are on or not :/

OpenStudy (jennyrlz):

So I will work it out with as much detail as I can provide :)

OpenStudy (jennyrlz):

First we solve the parenthesis (p) 8-3=5

OpenStudy (jennyrlz):

(5)^2 we still have a parenthesis

OpenStudy (anonymous):

p = (8-3) = 5 25 divided by 5 ^2 -1 e ^2 = 5 to te 2 power? so it'd be 25 25 divided by 25 - 1 m no multiplication right? :o double check >O< d 25 divided by 25 equals 1 :o a no addition i think right? sos it'd be 1 - 1 s 1-1= 0 o-o

OpenStudy (jennyrlz):

So then we sole the parenthesis 5^2 or 5x5 =25

OpenStudy (anonymous):

ohs sorry v-v i didn't see you replied @jennyrlz Q.Q

OpenStudy (jennyrlz):

No problem you saved me a lot of time xD

OpenStudy (jennyrlz):

You medal me I medal you everyone is happy :)

OpenStudy (anonymous):

-gives chu te medal anyways- :D

OpenStudy (anonymous):

awws <3

OpenStudy (anonymous):

hehe

OpenStudy (jennyrlz):

Well cya around :p

OpenStudy (anonymous):

Bye byee :P o-o my kik is Darkataevin ! NEW FRIEND~

OpenStudy (anonymous):

lol